Santos Unlimited (SU) was originally unlevered with 4800 shares outstanding. However, after a major financial restructure, SU now has $38000 of debt, with an annual interest expense of 12 percent. The restructuring has reduced the number of shares to 3000. A group of shareholders of SU are not convinced that this move towards adopting financial leverage is a good idea. Their main argument is that there is now some range of EBIT, however low, that will make the shareholders worse off than before. Help understand the situation better by computing the level of earnings before interest and tax (EBIT) that would make shareholders indifferent between being unlevered (i.e. not having any debt) and levered (i.e. having debt). Assume a 32 percent corporate tax rate. Answer: $ Place your answer to the nearest dollar without a dollar sign or a comma (if applicable).
Answer #1

Let the level of EBIT be x
x*(1-32%)/4800=(x-38000*12%)*(1-32%)/3000
=>3x/8=38000*12%
=>x=38000*12%*8/3
=>x=12160